OpenStudy (booklover40):

If a,b,c and d are replaced by four distinct digits from 1 to 9, then what's the largest possible value of the difference a.b - c.d? Help!!!!!

OpenStudy (booklover40):

Anyone? Help!

OpenStudy (booklover40):

Nobody?

OpenStudy (booklover40):

This isn't making day better. :(

OpenStudy (booklover40):

Nobody has even commented, and this has been up 20 minutes.

OpenStudy (booklover40):

Nobody has even commented, and this has been up 20 minutes.

OpenStudy (anonymous):

a=9 b=8 c=1 d=2 ab-cd=9*8-1*2=72-2=70

OpenStudy (anonymous):

value of a and b can interchange. similarly value of c and d can interchange.